OpenStudy (anonymous):

Solve for the variable. (t – 3)7 = 6t + 21 t = ____ Enter a numerical answer.

OpenStudy (anonymous):

t= 42

OpenStudy (anonymous):

(t-3)7=6t+21 7t-21=6t+21 7t=6t+42 t=42 *if you confused :D